20. Solve: 5^(2x-3) = 6^3x.



Answer by stanbon(48535)

Solve: 5^(2x-3) = 6^3x.
Take the log of both sides to get:
(2x-3)log5 = 3xlog6
Get the x terms together on one side:
Solve: 5^(2x-3) = 6^3x.
[2log5-3log6]x = 3log5
x = 3log5/[2log5-3log6]
x = 2.0969.../0.85157...
x = 2.4624..
Cheers,
Stan H.